OpenStudy (anonymous):

AB is a chord of a circle having center O , if angle AOB =90degree, prove that the chord AB is of length r root 2 , where r is the radius of circle

OpenStudy (anonymous):



OpenStudy (anonymous):

USE pythagorean theorem in triangle AOB

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

AB^2 =r^2 +r^2 AB^2=2r^2 AB=r root2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

i think should we need to draw a perpendicular